### v1

# Edge-guided Low-light Image Enhancement with Inertial Bregman Alternating Linearized Minimization

Abstract
Prior-based methods for low-light image enhancement often face challenges in extracting available prior information from dim images. To overcome this limitation, we introduce a simple yet effective Retinex model with the proposed edge extraction prior. More specifically, we design an edge extraction network to capture the fine edge features from the low-light image directly. Building upon the Retinex theory, we decompose the low-light image into its illumination and reflectance components and introduce an edge-guided Retinex model for enhancing low-light images. To solve the proposed model, we propose a novel inertial Bregman alternating linearized minimization algorithm. This algorithm addresses the optimization problem associated with the edge-guided Retinex model, enabling effective enhancement of low-light images. Through rigorous theoretical analysis, we establish the convergence properties of the algorithm. Besides, we prove that the proposed algorithm converges to a stationary point of the problem through nonconvex optimization theory. Furthermore, extensive experiments are conducted on multiple real-world low-light image datasets to demonstrate the efficiency and superiority of the proposed scheme.
Index Terms:
Edge-prior, Retinex, image enhancement, alternating linearized minimization, inertial, convergence analysis
I Introduction
Due to lighting and environmental conditions, images captured by the camera may be low-light, which negatively impacts post-processing tasks like segmentation, classification, and recognition. Therefore, it is crucial to improve the quality of low-light images. Note that the Retinex theory describes the low-light image as the reflectance and the illustration part [1, 2]. By enhancing the reflectance and adjusting the illustration at the same time, the norm-light image can be obtained [3, 4]. Traditionally, the prior information is used to help enhance the reflectance [5, 6]. However, the detailed information, such as edge information, contained in the low-light image is difficult to obtain directly. As shown in Fig. 1, observed image (a) with low contrast, one can hardly recognize there is a bicycle, which is clear in the enhanced image (e). Besides, the edge information is hard to capture entirely by the traditional method as shown in (b). Hence, resulting the unsatisfied results (d). In this paper, we propose to learn clear edge information directly from the observed dim image to enhance the quality of the low-light image.
(a) Input
(b) Edge
(c) Our edge
(d) Retinex with (b)
(e) Ours
Figure 1: Low-light image enhancement. (a) is the observed low-light image; (b) is the edge obtained from (a) by gradient and divergence operator; (c) is the edge obtained from (a) by the proposed scheme; (d) is the result of Retinex method with edge (b); (e) is the result of our method with our edge (c). We evaluate the enhanced results with the Google Vision API and the probability of our results has a Bicycle being judged significantly.
Generally, a low-light image $S$ can be decomposed as the reflectance part $R$ and illumination part $L$ [7], i.e., $S=R\circ L$, where $R\in(0,1)$ and $L\in(0,\infty)$. Directly enhancing the corresponding $R$ and $L$ to reconstruct the desired image $\hat{S}$ may be challenged for both $R$ and $L$ are unknown. Hence a minimization model was proposed as
$\min_{R,L}\frac{\beta}{2}\|S-R\circ L\|^{2}+\frac{\alpha}{2}\Psi(L)+\Phi(R),$ (1)
where the first term is the data fidelity term, $\Phi(\cdot)$ and $\Psi(\cdot)$ are the prior regularizers for $R$ and $L$, respectively, $\beta$ and $\alpha$ are positive parameters to balance the different terms. Based on the Retinex model, many crafted prior were designed to tackle the low-light image enhancement task. To better preserve fine details, Gu et al. [8] utilized the modified total variational model to fractional order for low-light image enhancement. The ADMM with convergence guarantee was proposed by adding auxiliary variables to solve the model. However, when their model faces the non-uniformed light image, the enhanced image may be unsatisfied. To overcome this disadvantage, Ma et al. [9] improved the model with a weighted map without theoretical analysis. By extending the fractional-order gradient total variation, small-magnitude details are maintained in the reflectance and the illumination is more likely piecewise smooth.
However, lower contrast often indicates smaller gradient magnitudes. Accordingly, the edge information cannot be obtained directly from the observed image. Moreover, the computational complexity of these models is expensive for the product form of the unknown variables. To tackle this problem, the logarithmic domain can be considered. Let $s=\log S$, $l=\log L$, and $r^{\prime}=\log R$, then we have $s=l+r^{\prime}$. Note that $R\in(0,1)$, we let $r=-r^{\prime}>0$, then $l=s+r$. Hence Eq. (1) can be reformulated as
$\min_{r,l}\frac{\beta}{2}\|l-s-r\|^{2}+\frac{\alpha}{2}\Psi(l)+\Phi(r).$ (2)
Based on the model in the logarithmic domain, a total variation model was proposed in [10]. They let $\Psi(l)=\|\nabla l\|^{2}+\frac{\mu}{2}\|l\|^{2}$ and $\Psi(r)=\|\nabla r\|_{1}$ to analyze the convergence of the proposed image enhancement model. Later, a dictionary learning-based model was proposed in [11] to improve Eq. (2) by saving more details through sparse representation. While discussing the Retinex model in the logarithmic domain, the edge information still cannot be fully used. Thus, how to get a more clear edge from the low-light image becomes a bottleneck.
Methods Fide. Reg. Learn. Alg. Conv.
[8]
[9]
[10]
[11]
[12]
[13]
[14]
[15]
[16]
[17]
[18]
Ours
  • 1
    ‘Fide.’ means that the data fidelity term holds; ‘Reg.’ represents preserving the regularizer; ‘Learn.’ denotes involving the deep learning-based prior; ‘Alg.’ implies proposing the new algorithm; ‘Conv.’ indicates analyzing the convergence of the algorithm.
TABLE I: SUMMARIZING OF ESSENTIAL COMPARISON ASPECTS OF Retinex-based METHODS1.
In [12], the authors attempted to manipulate the gradient magnitudes of the reflectance to boost the contrast. By multiplying the gradient of the reflectance part with a constant, a more clear edge information can be collected. Similarly, a dual reflectance estimation method was proposed in [13] with convergence analysis to improve the performance of the Retinex model. In [14], a weighted map was designed to adjust the reflectance. Moreover, an alternating proximal gradient algorithm was proposed with convergence analysis to solve their Retinex model. Following this line, the weighted fractional-order total variation and dark channel prior were applied in [19] to amplify the edge information. Inspiring the hybrid deep learning prior [20, 21], Wu et al. [15] improved this model by adding a denoiser to remove the noise in the image. Instead of learning the data-fitting directly as RetinexNet [16], Liu et al. [17] further replaced $\Psi(L)$ with two implicit regularizers. A similar idea was presented in [18]. These works have competitive results in enhancing low-light images, however, their model and algorithm lack the theoretical guarantee. Besides, from these outstanding works, we know that edge information is essential to the low-light image enhancement task.
As mentioned before, extracting the edge information directly from the observed low-light image turns into unsatisfying results. Note that deep learning-based methods have revealed prominent performance in image processing. We wondering if the deep learning-based methods can help us to improve the results of the variational Retinex model. Typically, the edge $G$ of an image $S$ can be found by
$G=\mbox{div}(S_{x},S_{y}),$ (3)
where $S_{x}=\frac{\nabla_{x}S}{\sqrt{1+\left|\nabla S\right|^{2}}}$, $S_{y}=\frac{\nabla_{y}S}{\sqrt{1+\left|\nabla S\right|^{2}}}$, and $\mbox{div}(\cdot)$ denotes the divergence operation. However, when it comes to low-light images, as shown in Fig. 1 (b), the edge information is unsatisfied. Note that the authors in [22] applied the deep learning-based module to generate an edge prediction network to help the image super-resolution task. With such edge prediction network, the edge information of the low-resolution image can be extracted easily. A similar idea was used in [23] and [24] for image denoising and in [25] for MRI reconstruction.
Inspired by these works, we aim to explore the deep edge prediction network to extract the edge feature from the low-light image and integrate the learned edge prior to the Retinex model. For simplicity, we will focus on the Retinex model in the logarithmic domain. To solve our low-light image enhancement model, we propose a novel inertial Bregman alternating linearized minimization algorithm. The details of the proposed algorithm will be introduced in Section II-C. For a clear perception, we summarize the essential comparison aspects in Table I. The contributions of the proposed low-light image enhancement scheme can be summarized as follows.
  • We propose an edge-guided prior for low-light image enhancement, where the latent fine edge information can be exacted directly from the observed low-light image.
  • With the help of the latent edge information, we integrate the proposed edge-guided prior to the Retinex model in the logarithmic domain for low-light image enhancement.
  • To solve the proposed model, an effective inertial Bregman alternating linearized minimization algorithm is proposed. Besides, we establish the convergence properties of the proposed algorithm.
  • To illustrate the robustness and superiority of the proposed scheme over several state-of-the-art methods, experiments in real-world low-light data are conducted.
The remainder of this paper is organized as follows. Section II reviews the basic concept of the Retinex theory, the edge prior, and some preliminaries for the proposed algorithm. The proposed edge extraction, image enhancement model, and algorithm are displayed in Section III. In Section IV, the convergence properties of the proposed algorithm are established. Experimental results on different datasets are in Section V. Finally, the conclusions follow in Section VI.
II Related Works
In this section, we will give a short recall of Retinex theory on image processing, the edge prior, and some preliminaries for the proposed algorithm.
II-A Retinex image enhancement
The goal of Retinex theory [1] is to describe how the human visual system processes color changes and interprets the color of objects/scenes under changing illumination circumstances. In other words, Retinex theory explains that an object will show different colors under different lights, but people can know what color the object should be through common sense and visual senses. Therefore, the observed $S$ under different light can be described as
$S=R\circ L,$ (4)
where $R$ is the reflectance of the surface, and $L$ is the illumination on the surface.
With the product of reflectance and illuminance of an image, the Retinex-based image processing system is constructed. Hence, we only need to find $R$ and $L$ of a low-light image and enhance the intensity. However, it is difficult to find $R$ and $L$, directly. Consequently, model (1) is proposed with different regularizers to find better $R$ and $L$. Meanwhile, the logarithmic domain model (2) is proposed to reduce the computation complexity. In this paper, we focus on the model in the logarithmic domain for low-light image enhancement.
II-B Edge prior
If we find the edge of an image, the most detailed information of the image can be restored [26]. The variational method to extract the edge of an image $S$ is to constrain the gradient $\nabla S$. The discrete gradient operator $\nabla:\mathbb{R}^{m\times n}\rightarrow\left(\mathbb{R}^{m\times n},\ \mathbb{R}^{m\times n}\right)$ is defined by
$\vspace{-0.15in}(\nabla S)_{i,j}=\left(\left(\nabla_{x}S\right)_{i,j},\left(\nabla_{y}S\right)_{i,j}\right)$
with
$\left(\nabla_{x}S\right)_{i,j}=\left\{\begin{array}[]{lll}S_{i+1,j}-S_{i,j},&\text{ if }&i<m\\ S_{1,j}-S_{m,j},&\text{ if }&i=m\end{array}\right.$
and
$\left(\nabla_{y}S\right)_{i,j}=\left\{\begin{array}[]{lll}S_{i,j+1}-S_{i,j},&\text{ if }&j<n\\ S_{i,1}-S_{i,n},&\text{ if }&j=n\end{array}\right.$
for $i=1,\ldots,m$ and $j=1,\ldots,n$, where $S_{i,j}$ represents the value of pixel $(i,j)$ in the image.
Usually, the total variation model (ROF) [27] was proposed to preserve the edge of an image as
$\|\nabla S\|_{1}=\sum_{i,j}\left|(\nabla S)_{i,j}\right|=\sum_{i,j}\sqrt{\left(\nabla_{x}S\right)_{i,j}^{2}+\left(\nabla_{y}S\right)_{i,j}^{2}}.$
There are many variants of the ROF model to represent the edge of the image, such as the change of norm, weighted gradient, generalized variation, ROF model in different spaces, etc. More specifically, the ROF model was used in [28] as an additional regular term to obtain the detailed texture in hyperspectral anomaly detection. The ROF model with $\ell_{0}$-norm $\|\nabla S\|_{0}$ was utilized to enforce the sparsity of image gradient to help the image restoration in [29]. More structural information was collected by applying a weighted ROF regularizer $\|T\nabla S\|_{1}$ for image segmentation [30]. Total generalized variation (TGV) contributed to help maintain sharp edges in [31]. Total variation in HSV color space was applied to help image enhancement [32]. Nonlocal total variation was applied to data speckle reduction in [33].
From the above literature, the edge of the image has its advantages for different tasks. In this paper, we learn the edge by using the convolutional neural network to help the image enhancement task.
Figure 2: The flowchart of the proposed scheme. We use 485 paired images from the LOL dataset as the training sets. The edge information can be directly obtained by using the proposed Edge-Net. The yellow, green, and blue blocks are the convolution layer, multi-scale residual block, and bottleneck layer, respectively. Then we use the learned edge prior to helping the image Retinex task. Through the proposed convergent algorithm, the low-light image can be well enhanced.
II-C Preliminaries
In this subsection, we review some preliminaries about subdifferential and Kurdyka-Łojasiewicz (KL) property in nonconvex optimization, which will be used in our algorithm design and convergence analysis.
Definition 1.
[34, 35] (Subdifferentials) Let $f:\mathbb{R}^{n}\rightarrow(-\infty,+\infty]$ be a proper and lower semicontinuous function.
  • (i)
    For a given $x\in{\rm dom}f$, the Fréchet subdifferential of $f$ at $x$, written by ^$\widehat{\partial}f(x)$, is the set of all vectors $u\in\mathbb{R}^{n}$ satisfying
    $\liminf_{y\neq x,y\rightarrow x}\frac{f(y)-f(x)-\langle u,y-x\rangle}{\|y-x\|}\geq 0,$
    and we set ^$\widehat{\partial}f(x)=\emptyset$ when $x\notin{\rm dom}f$.
  • (ii)
    The limiting-subdifferential, or simply the subdifferential, of $f$ at $x$, written by $\partial f(x)$, is defined by
    $\partial f(x):=\{u\in\mathbb{R}^{n}\;|\;\exists\ x^{k}\rightarrow x,~{}{\rm s.t.}~{}f(x^{k})\rightarrow f(x)$ (5)
    $~{}{\rm and}~{}\widehat{\partial}f(x^{k})\ni u^{k}\rightarrow u\}.$
  • (iii)
    A point $x^{*}$ is called (limiting-)critical point or stationary point of $f$ if it satisfies $0\in\partial f(x^{*})$, and the set of critical points of $f$ is denoted by ${\rm crit}f$.
We use the notation ${\rm dom}(\partial f):=\{x\in\mathbb{R}^{n}\;|\;\partial f(x)\neq\emptyset\}$. If $g$ is a continuously differentiable function, it holds that $\partial(f+g)=\partial f+\nabla g$.
Next, we recall the KL property [36, 35, 37], which plays a central role in the convergence analysis.
Definition 2.
(KL property and KL function) Let $f:\mathbb{R}^{n}\rightarrow(-\infty,+\infty]$ be a proper and lower semicontinuous function.
  • $(a)$
    $f$ is said to satisfy KL property at $x^{*}\in{\rm dom}(\partial f)$ if there exist a constant $\eta\in(0,+\infty]$, a continuous concave function $\phi:[0,\eta)\rightarrow\mathbb{R}_{+}$ and a neighborhood $U$ of $x^{*}$, such that
    • (i)
      $\phi(0)=0$ and $\phi$ is continuously differentiable on $(0,\eta)$ with $\phi^{\prime}>0$;
    • (ii)
      for all $x\in U\cap\{z\in\mathbb{R}^{n}\;|\;f(x^{*})<f(z)<f(x^{*})+\eta\}$, the following inequality holds:
      $\phi^{\prime}(f(x)-f(x^{*})){\rm dist}(0,\partial f(x))\geq 1.$ (6)
  • $(b)$
    $f$ is called a KL function if $f$ satisfies the KL property at each point of dom$(\partial f)$.
Denote $\Phi_{\eta}$ the set of functions $\phi$ which satisfy the involved conditions in Definition 2(a).
Definition 3.
(Bregman distance) Let $\psi(\cdot):\mathbb{R}^{n}\rightarrow(-\infty,+\infty]$ be a strongly convex and continuously differentiable function over $\rm{dom}(\partial\psi)$. The Bregman distance associated with the kernel $\psi(\cdot)$ is the function $D_{\psi}(\cdot,\cdot):\rm{dom}(\psi)\times\rm{dom}(\partial\psi)\rightarrow\mathbb{R}$ given by
$D_{\psi}(x,y)=\psi(x)-\psi(y)-\langle\nabla\psi(y),x-y\rangle,$
$\qquad\qquad\forall x\in\operatorname{dom}(\psi),y\in\operatorname{int}(\operatorname{dom}(\psi)).$
Note that the Bregman distance includes the standard Euclidean distance as its special case, i.e., $D_{\psi}(x,y)=\frac{1}{2}\|x-y\|^{2}$ if the kernel function is chosen by $\psi(\cdot)=\frac{1}{2}\|\cdot\|^{2}$. We refer to [38, 39, 40, 41, 42] and references therein for the Bregman distance.
III Edge-guided model with algorithm
III-A Edge extraction
Different from the traditional edge extraction methods, in this paper, we propose to obtain the edge from a simple but efficiently designed neural network. More specifically, we use the multi-scale residual block (MSRB) [43] as the backbone of the proposed edge network. Due to the bypass network structure, the MERB can detect more features of images and the information of different passes can be shared with each other. Giving the observed image $S$ and the corresponding label $\widehat{S}$, the proposed edge detection network is trained. The loss function of the proposed network is
$\mathcal{L}=\|G(S)-\widehat{S}_{egde}\|_{1},$ (7)
where $G(\cdot)$ represents the proposed edge network, $G(S)$ denotes the trained edge, and $\widehat{S}_{egde}$ is the edge label. Here the truth edge label is obtained from Eq. (3). As shown in Fig.2, we capture labels from the corresponding bright-light images. By minimizing the loss function, we train the edge network that can generate the edge directly from the low-light image. The detailed training setting is described in Section V.
III-B Retinex image enhancement model
As mentioned before, the intensity of $S$ mainly depends on the amount of source illumination falling on the objects in the scene and the amount of illumination reflected by the objects in the scene. Based on these two components, we can introduce two functions: the illumination function $L$ and the reflection function $R$. These two functions are assumed to be combined as a product to form image $S$
$S=L\circ R.$ (8)
Due to the reflectivity, $R\in(0,1)$. For the illumination effect, $L\in(0,\infty)$. Then, we know that $L>S>0$. To tackle the product form, we convert it into the logarithmic domain following [10], i.e.,
$s=\log(S),\quad l=\log(L),\quad r^{\prime}=\log(R).$ (9)
Then we obtain
$s=l+r^{\prime}.$ (10)
Note that $10<R<$ 1, hence $r^{\prime}<0$. We set $r=-r^{\prime}>0$, then the above model becomes
$l=s+r.$ (11)
Based on the former assumption, we propose the following energy functional for Retinex
$E(r,l)=\|\nabla l\|_{1}+\frac{\alpha}{2}\|\nabla r-\nabla g\|^{2}+\frac{\beta}{2}\|l-s-r\|^{2},$ (12)
where $g=G(S)$ is the learned edge prior, $\alpha$ and $\beta$ are the positive parameters.
III-C Inertial Bregman Alternating Minimization Algorithm
To solve the proposed model (12), we consider a more general class of structured optimization problems that minimize the following function:
$\Phi(x_{1},x_{2}):=\theta_{1}(x_{1})+\underbrace{h^{+}(x_{1},x_{2})-h^{-}(x_{1},x_{2})}_{h(x_{1},x_{2})}+\theta_{2}(x_{2}),$ (13)
where $x_{1}$ and $x_{2}$ are decision variables, $\theta_{1}$ and $\theta_{2}$ are assumed to be general nonsmooth and nonconvex functions, and $h(x_{1},x_{2})$ is continuously differentiable for both $x_{1}$ and $x_{2}$ component-wise, which is also assumed to be decomposable into two parts $h^{+}(x_{1},x_{2})$ and $h^{-}(x_{1},x_{2})$. Define
$\mathcal{P}_{g,\mu}^{\psi}(u;v):=\arg\min_{\hskip-11.38092ptx\in\mathbb{R}^{n}}\left\{g(x)+\langle x,v\rangle+\mu D_{\psi}(x,u)\right\}$ (14)
with certain strongly convex function $\psi$ and the Bregman divergence $D_{\psi}(\cdot,\cdot)$.
In the following, we propose an inertial Bregman alternating linearized minimization algorithm to solve the general optimization problem that minimizes the function given in (13). The algorithm framework is summarized in Algorithm 1. In Algorithm 1, the kernel functions $\psi_{k}$ and $\varphi_{k}$, for all $k$, are selected to be strongly convex with the smallest moduli $\rho_{1}$ and $\rho_{2}$, respectively. The most related works are [44, 45, 46, 41, 47].
Algorithm 1 Inertial Bregman alternating linearized minimization algorithm
  Choose the parameters $\alpha_{i}^{k}$, $\beta_{i}^{k}$ and $\tau_{i}^{k}$, $i=1,2$. For given $x^{0}\in\mathbb{R}^{n}$, $x^{-1}:=x^{0}$, and $k:=0$.
  while the stopping criterion is not satisfied, do
      1. Compute $y_{1}^{k},z_{1}^{k}$ and $x_{1}^{k+1}$ by
$y_{1}^{k}:=x_{1}^{k}+\alpha_{1}^{k}(x_{1}^{k}-x_{1}^{k-1}),$ (15a)
$z_{1}^{k}:=x_{1}^{k}+\beta_{1}^{k}(x_{1}^{k}-x_{1}^{k-1}),$ (15b)
$x_{1}^{k+1}\in{\bf\mathcal{P}}_{\theta_{1}+h^{+}(\cdot,x_{2}^{k}),\tau_{1}^{k}}^{\psi_{k}}\left(x_{1}^{k};t_{1}^{k}\right),$ (15c)
with $t_{1}^{k}=-\nabla h^{-}(z_{1}^{k},x_{2}^{k})-\tau_{1}^{k}(y_{1}^{k}-x_{1}^{k})$. 2. Compute $y_{2}^{k},z_{2}^{k}$ and $x_{2}^{k+1}$ by
$y_{2}^{k}:=x_{2}^{k}+\alpha_{2}^{k}(x_{2}^{k}-x_{2}^{k-1}),$ (16a)
$z_{2}^{k}:=x_{2}^{k}+\beta_{2}^{k}(x_{2}^{k}-x_{2}^{k-1}),$ (16b)
$x_{2}^{k+1}\in{\bf\mathcal{P}}_{\theta_{2}+h^{+}(x_{1}^{k+1},\cdot),\tau_{2}^{k}}^{\varphi_{k}}\left(x_{2}^{k};t_{2}^{k}\right),$ (16c)
with $t_{2}^{k}=-\nabla h^{-}(x_{1}^{k+1},z_{2}^{k})-\tau_{2}^{k}(y_{2}^{k}-x_{2}^{k})$.
  end while
Next, we apply the proposed inertial Bregman alternating minimization algorithm, i.e., Algorithm 1, to solve (12). Note that the model (12) falls into the form of (13) with $(x_{1},x_{2}):=(l,r)$, $\theta_{1}(x_{1}):=\|\nabla l\|_{1}$, $\theta_{2}(x_{2}):=\frac{\alpha}{2}\|\nabla r-\nabla g\|^{2}$, $h^{+}(x_{1},x_{2}):=0$, and $h^{-}(x_{1},x_{2}):=-\frac{\beta}{2}\|l-s-r\|^{2}$. By choosing $\psi_{k}\equiv\psi:=\frac{1}{2}\|\cdot\|^{2}$ and $\varphi_{k}:=\frac{1}{2}\|\cdot\|_{M_{k}}^{2}$ with $M_{k}:=\gamma_{k}I-\frac{\alpha}{\tau_{2}^{k}}\nabla^{\top}\nabla$, where $\gamma_{k}>\frac{\alpha}{\tau_{2}^{k}}\lambda_{\max}(\nabla^{\top}\nabla)$ and applying Algorithm 1 to solve (12), we obtain the following $l$- and $r$-subproblems:
[…]
$\qquad\qquad+\langle l,\beta(r^{k}-z_{1}^{k}-s)-\tau_{1}^{k}(y_{1}^{k}-l^{k})\rangle\Big{\}},$
and
[…]
$\qquad\qquad+\langle r,\beta(l^{k+1}-z_{2}^{k}+s)-\tau_{2}^{k}(y_{2}^{k}-r^{k})\rangle\Big{\}}.$
Notice that the $l$-subproblem (17) is a convex optimization problem that can be efficiently solved by classic ADMM algorithm, while the $r$-suproblem (18) possesses the closed-form solution according to the definition of $M_{k}$ as follows:
$r^{k+1}=\frac{\alpha\nabla^{T}\nabla g+\beta(z_{2}^{k}-l^{k+1}-s)+\tau_{2}^{k}(y_{2}^{k}+M_{k}r^{k}-r^{k})}{\gamma^{k}}.$ (19)
Therefore, the detailed iterative scheme of Algorithm 1 for solving the proposed model can be read as follows:
[…]
In the implementation, the parameter $\gamma_{k}$ is not difficult to choose since the maximum eigenvalue $\lambda_{\max}(\nabla^{\top}\nabla)\leq 8$ [48].
IV Convergence results
Now we are prepared to establish the convergence of the proposed Algorithm 1. We start by giving some standard assumptions for (13).
Assumption 1.
The components $\theta_{i}(i=1,2)$ are proper lower semicontinuous functions. The objective function $\Phi(x_{1},x_{2})$ is bounded from below for all $x_{1},x_{2}$.
Assumption 2.
The function $h(x_{1},x_{2})=h^{+}(x_{1},x_{2})-h^{-}(x_{1},x_{2})$ satisfies the following conditions.
  • (i)
    For any fixed $x_{2}$, the function $h^{-}(\cdot,x_{2})$ is $C_{L_{1}^{-}(x_{2})}^{1,1}$, i.e., the partial gradient $\nabla_{x_{1}}h^{-}(\cdot,x_{2})$ is globally Lipschitz with module $L_{1}^{-}(x_{2})$, that is for $\forall u,v\in\mathbb{R}^{n}$,
    $\left\|\nabla_{x_{1}}h^{-}\left(u,x_{2}\right)-\nabla_{x_{1}}h^{-}\left(v,x_{2}\right)\right\|\leq L_{1}^{-}(x_{2})\left\|u-v\right\|.$
    Likewise, for any fixed $x_{1}$, the function $h^{-}(x_{1},\cdot)$ is assumed to be $C_{L_{2}^{-}(x_{1})}^{1,1}$.
  • (ii)
    For $i=1,2$, there exist $\lambda_{i}^{-}$ and $\lambda_{i}^{+}>0$ such that $\inf\left\{L_{1}^{-}\left(x_{2}^{k}\right):k\in\mathbb{N}\right\}\geq\lambda_{1}^{-}$, $\inf\left\{L_{2}^{-}\left(x_{1}^{k}\right):k\in\mathbb{N}\right\}\geq\lambda_{2}^{-}$, $\sup\left\{L_{1}^{-}\left(x_{2}^{k}\right):k\in\mathbb{N}\right\}\leq\lambda_{1}^{+}$, and $\sup\left\{L_{2}^{-}\left(x_{1}^{k}\right):k\in\mathbb{N}\right\}\leq\lambda_{2}^{+}.$
  • (iii)
    $h$ is a continuously differentiable function and its gradient $\nabla h(x_{1},x_{2})=\nabla h^{+}(x_{1},x_{2})-\nabla h^{-}(x_{1},x_{2})$ is Lipschitz continuous on bounded subsets of $\mathbb{R}^{n}\times\mathbb{R}^{m}$. That is, for each bounded subsets $\mathbb{B}_{1}\times\mathbb{B}_{2}$ of $\mathbb{R}^{n}\times\mathbb{R}^{m}$, there exists $L>0$ (or $L^{+}$and $L^{-}$satisfying $L=L^{+}+L^{-}$) such that
    $\left\|\left(\begin{array}[]{l}\nabla_{x_{1}}h\left(x_{1},y_{1}\right)\\ \nabla_{x_{2}}h\left(x_{1},y_{1}\right)\end{array}\right)-\left(\begin{array}[]{c}\nabla_{x_{1}}h\left(x_{2},y_{2}\right)\\ \nabla_{x_{2}}h\left(x_{2},y_{2}\right)\end{array}\right)\right\|$
    $\leq L\left\|\left(\begin{array}[]{l}x_{1}\\ y_{1}\end{array}\right)-\left(\begin{array}[]{l}x_{2}\\ y_{2}\end{array}\right)\right\|$
    for all $\left(x_{i},y_{i}\right)\in\mathbb{B}_{1}\times\mathbb{B}_{2}~{}(i=1,2)$.
Assumption 3.
The parameters $\{\alpha_{i}^{k}\}$, $\{\beta_{i}^{k}\}$, and $\{\tau_{i}^{k}\}$, $i=1,2$, for all $k\in\mathbb{N}$ should satisfy the following conditions:
  • (i)
    there exist $0<\bar{\alpha}_{i}<\frac{(1-\varepsilon)\rho_{i}}{2}$ such that $\alpha_{i}^{k}\in[0,\bar{\alpha}_{i}]$, $i=1,2$. In addition, $\beta_{i}^{k}\in[0,\bar{\beta}_{i}]$ for some $\bar{\beta}_{i}>0$, $i=1,2$.
  • (ii)
    it holds that
    $\tau_{1}^{k}=\frac{(1+\varepsilon)\delta_{1}+(1+\beta_{1}^{k})L_{1}^{-}(x_{2}^{k})}{\rho_{1}-\alpha_{1}^{k}},$ (21)
    $\tau_{2}^{k}=\frac{(1+\varepsilon)\delta_{2}+(1+\beta_{2}^{k})L_{2}^{-}(x_{1}^{k+1})}{\rho_{2}-\alpha_{2}^{k}},$
    where
    $\delta_{1}=\frac{\bar{\alpha}_{1}+\bar{\beta}_{1}\rho_{1}}{(1-\varepsilon)\rho_{1}-2\bar{\alpha}_{1}}\lambda_{1}^{+},~{}~{}\delta_{2}=\frac{\bar{\alpha}_{2}+\bar{\beta}_{2}\rho_{2}}{(1-\varepsilon)\rho_{2}-2\bar{\alpha}_{2}}\lambda_{2}^{+}.$ (22)
Remark 1.
Note that according to Assumption 3, for all $k\in\mathbb{N}$, $\tau_{i}^{k},i=1,2$ is bounded from above, that is, $\tau_{i}^{k}\leq\tau_{i}^{+},i=1,2,$ with
$\tau_{1}^{+}=\frac{(1+\varepsilon)\delta_{1}+(1+\bar{\beta}_{1})\lambda_{1}^{+}}{\rho_{1}-\bar{\alpha}_{1}},$ (23)
$\tau_{2}^{+}=\frac{(1+\varepsilon)\delta_{2}+(1+\bar{\beta}_{2})\lambda_{2}^{+}}{\rho_{2}-\bar{\alpha}_{2}}.$
Before the convergence analysis, for notational simplicity, we define an auxiliary function that enjoys the property of yielding a sequence of decreasing function values. Let $\Theta:\mathbb{R}^{n_{1}\times n_{2}}\times\mathbb{R}^{n_{1}\times n_{2}}\rightarrow(-\infty,\infty]$ be the auxiliary function
$\Theta_{\delta_{1},\delta_{2}}(\mathbf{w}):=\Phi\left(w_{1}\right)+\frac{\delta_{1}}{2}\left\|w_{11}-w_{21}\right\|^{2}+\frac{\delta_{2}}{2}\left\|w_{12}-w_{22}\right\|^{2},$ (24)
where $\Phi$ is defined in (13), $\delta_{1},\delta_{2}$ are defined in (22), $w_{1}=\left(w_{11},w_{12}\right)\in\mathbb{R}^{n_{1}}\times\mathbb{R}^{n_{2}},w_{2}=\left(w_{21},w_{22}\right)\in\mathbb{R}^{n_{1}}\times\mathbb{R}^{n_{2}}$, and $\mathbf{w}=\left(w_{1},w_{2}\right)$.
Lemma 1.
Let $\left\{\left(x_{1}^{k},x_{2}^{k}\right)\right\}_{k\in\mathbb{N}}$ be a sequence generated by (1), for all $k\in\mathbb{N}$, $w_{1}^{k}=$ $\left(x_{1}^{k},x_{2}^{k}\right),w_{2}^{k}=\left(x_{1}^{k-1},x_{2}^{k-1}\right)$, and $\mathbf{w}^{k}=\left(w_{1}^{k},w_{2}^{k}\right)$. Suppose that Assumptions 1, 2 and 3 hold. Then, the sequence $\{\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k})\}_{k\in\mathbb{N}}$ is nonincreasing. In particular, for all $k\in\mathbb{N}$ and $\varepsilon>0$, it holds that
$\Theta_{\delta_{1},\delta_{2}}\left(\mathbf{w}^{k+1}\right)-\Theta_{\delta_{1},\delta_{2}}\left(\mathbf{w}^{k}\right)\leq-\delta\left\|\mathbf{w}^{k+1}-\mathbf{w}^{k}\right\|^{2},$ (25)
where $\delta_{1},~{}\delta_{2}$ are defined in (22), and $\delta=\frac{\varepsilon}{2}\min\{\delta_{1},\delta_{2}\}>0$.
The proof of Lemma 1 can be found in the supplementary material Appendix A.
Lemma 2.
Let $\left\{\left(x_{1}^{k},x_{2}^{k}\right)\right\}_{k\in\mathbb{N}}$ be a sequence generated by (1), for all $k\in\mathbb{N}$, $w_{1}^{k}=$ $\left(x_{1}^{k},x_{2}^{k}\right),w_{2}^{k}=\left(x_{1}^{k-1},x_{2}^{k-1}\right)$, and $\mathbf{w}^{k}=\left(w_{1}^{k},w_{2}^{k}\right)$. Suppose that Assumptions 1, 2 and 3 hold. Then, for any $k\geq 1$, there exists $\gamma>0$ such that
${\rm dist}\left(0,\partial\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k+1})\right)\leq\gamma\left\|{\bf w}^{k+1}-{\bf w}^{k}\right\|.$ (26)
The proof of Lemma 2 can be found in the supplementary material Appendix B.
Theorem 1.
Let $\left\{\left(x_{1}^{k},x_{2}^{k}\right)\right\}_{k\in\mathbb{N}}$ be a sequence generated by (1), for all $k\in\mathbb{N}$, $w_{1}^{k}=$ $\left(x_{1}^{k},x_{2}^{k}\right),w_{2}^{k}=\left(x_{1}^{k-1},x_{2}^{k-1}\right)$, and $\mathbf{w}^{k}=\left(w_{1}^{k},w_{2}^{k}\right)$. Suppose that Assumptions 1, 2 and 3 hold, and $\{\mathbf{w}^{k}\}$ is assumed to be bounded. Then,
  • (i)
    it holds that $\sum_{k=0}^{\infty}\left\|\mathbf{w}^{k+1}-\mathbf{w}^{k}\right\|^{2}<\infty$, and furthermore $\lim_{k\rightarrow\infty}\left\|\mathbf{w}^{k+1}-\mathbf{w}^{k}\right\|=0.$
  • (ii)
    The set $z({\bf w}^{0})$ of the cluster point of $\left\{{\bf w}^{k}\right\}_{k\in\mathbb{N}}$ is nonempty and compact, and any point in $z({\bf w}^{0})$ is the critical point of (13).
  • (iii)
    The limit $\lim_{k\rightarrow\infty}\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k})=\Theta^{\infty}$ exist, and $\Theta_{\delta_{1},\delta_{2}}({\bf w}^{*})=\Theta^{\infty}$ for all ${\bf w}^{*}\in z({\bf w}^{0})$.
The proof of Theorem 1 can be found in the supplementary material Appendix C.
Theorem 2.
Let $\left\{\left(x_{1}^{k},x_{2}^{k}\right)\right\}_{k\in\mathbb{N}}$ be a sequence generated by (1), for all $k\in\mathbb{N}$, $w_{1}^{k}=$ $\left(x_{1}^{k},x_{2}^{k}\right),w_{2}^{k}=\left(x_{1}^{k-1},x_{2}^{k-1}\right)$, and $\mathbf{w}^{k}=\left(w_{1}^{k},w_{2}^{k}\right)$. Suppose that Assumptions 1, 2 and 3 hold, and $\{\mathbf{w}^{k}\}$ is assumed to be bounded. If $\Phi$ in (13) is a KL function. Then, the sequence $\left\{{\bf w}^{k}\right\}_{k\geq 1}$ has finite length, that is, $\sum_{k=1}^{\infty}\left\|{\bf w}^{k+1}-{\bf w}^{k}\right\|<+\infty.$ Hence, the whole sequence $\{{\bf w}^{k}\}_{k\geq 1}$ is convergent.
The proof of Theorem 2 can be found in the supplementary material Appendix D.
V Numerical experiments
To evaluate the effectiveness of the proposed scheme, we first analyze the proposed edge prior with some ablation study. Secondly, the parameters of the proposed model and algorithm are studied, respectively. Then, the real-world low-light images are tested to illustrated the robustness of our method. Furthermore, some state-of-the-art methods are compared to illustrate the superiority of the proposed approach.
Figure 3: Visual comparison of different edge extraction methods. The first column is low-light images; the second column and the last column are the edges obtained by using Eq. (3) and the learned edge by our strategy, respectively.
V-A Implement
In this work, we apply Eq. (3) on 485 training images of the LOL dataset [16] to obtain their corresponding edges. As illustrated in Fig. 2, we extract the edge information from the high-contrast image. Subsequently, we employ this edge feature as the label and train the edge network using the low-contrast input. The implementation of our edge network is carried out using the PyTorch framework, and we update it using the Adam optimizer. The initial learning rate is set to $10^{-4}$ and it is halved every 200 epochs. Once the edge network is learned, we integrate it into the proposed framework to enhance the performance of the image Retinex model.
Figure 4: Study on the number of multi-scale residual block (MSRB) modules. Considering the cost and the performance, we set the number of modules as 6 in our experiment.
Figure 5: Ablation study on edge prior. The images and the corresponding (color) histograms are presented. The first column is the input image, the second column is the result without edge prior, and the last column is the proposed result. Enhanced image with our edge prior has more clear detail.
V-B Effectiveness of edge prior
We propose an edge extraction network, which is comprised of the convolutional layer, multi-scale residual block (MSRB), upsample module, and bottleneck layer. To illustrate the effectiveness of the proposed edge network, Fig. 3 presents the visual comparison of the learned edges and the edges from Eq. (3). For the training, we study the effects of the number of MSRB modules. Theoretically, as the number of MSRB modules increases, the performance of the edge network gets better. In Fig. 4, the loss and PSNR curves show that the learned edges are closer to the labels with 6 MSRB modules. Considering the cost and the performance, in the experiments, we set the number of MSRB as 6 to capture the edge information directly from the input low-light images.
In Fig. 5, we compare the results of the proposed model with and without the edge prior. From the visual results of enhanced images in the first row, the results with the proposed learned edge prior show more detailed information. The second and the last rows display the histogram curves. For the histogram, the x-axis represents the variation in hue, while the y-axis represents the total number of pixels with that specific hue. In general, the left side of the x-axis corresponds to dark areas and the right side corresponds to light areas. Therefore, a histogram of a low-light image will predominantly exhibit data points on the left and center of the plot. In contrast, a histogram of an enhanced image will mostly show data points toward the right and center of the plot. In the last row, the red, green, and blue curves present the distribution of the red, green, and blue color channels, respectively. The histogram curves illustrate the competitive performance of the proposed scheme and the advantages of the proposed edge prior.
Figure 6: Comparison between the alternating minimization algorithm (AMA) and the proposed scheme.
V-C Algorithm and model analysis
In this subsection, we analyze the efficiency of the proposed scheme. First of all, we compare the proposed algorithm with the alternating minimization algorithm (AMA), which is referred to in literature [10]. The energy curves presented in Fig. 6 illustrate that with the same parameter setting, the proposed algorithm is faster and more accurate.
Secondly, the algorithm parameters $\alpha_{1}$, $\alpha_{2}$, $\beta_{1}$, $\beta_{2}$, and $\tau_{1}$, $\tau_{2}$ are studied. According to the Assumption 3, $\alpha_{i}^{k}\in[0,\bar{\alpha}_{i}]$ with $0<\bar{\alpha}_{i}<\frac{(1-\varepsilon)\rho_{i}}{2}$ and $\beta_{i}^{k}\in[0,\bar{\beta}_{i}]$ for some $\bar{\beta}_{i}\in(0,1)$, $i=1,2$. In experiments, we set strongly convex modulus $\rho_{1}$ and $\rho_{2}$ to be $1$, then we have $\alpha_{i}\in[0,0.5)$ and $\beta_{i}\in[0,1)$.
On the other hand, according to Assumption 2, $\lambda_{1}^{+}$ and $\lambda_{2}^{+}$ are set to be the supremum of $L_{1}^{-}(x_{2}^{k})$ and $L_{2}^{-}(x_{1}^{k})$, respectively. Note that $L_{1}^{-}\left(x_{2}^{k}\right)$ and $L_{2}^{-}\left(x_{1}^{k}\right)$ are the Lipschitz constants of $\nabla_{x_{1}}h^{-}\left(x_{1},x_{2}^{k}\right)$ at $x_{1}^{k}$ and $\nabla_{x_{2}}h^{-}\left(x_{1}^{k},x_{2}\right)$ at $x_{2}^{k}$, respectively.Since in the proposed model (12), $h^{-}(x_{1},x_{2})=\frac{\beta}{2}\|x_{2}-s-x_{1}\|^{2}$, we have $\|\nabla_{x_{1}}h^{-}(u,x_{2})-\nabla_{x_{1}}h^{-}(v,x_{2})\|=\beta\|u-v\|\leq L_{1}^{-}(x_{2})\|u-v\|$, hence $L_{1}^{-}(x_{2})\in[0,\beta]$. Similarly, we have $L_{2}^{-}(x_{1})\in[0,\beta]$. To this end, we know that for the proposed Retinex model (12), $\lambda_{1}^{+}$ and $\lambda_{2}^{+}$ are set to be $\beta$.
In the experiments, by setting $\bar{\beta}_{i}=1-\epsilon$, $\bar{\alpha}_{i}=\frac{1}{2}-\epsilon$, $\rho_{i}=1$, and $\lambda_{i}^{+}=\beta$, we get $\tau_{i}=\frac{\frac{1+\epsilon}{\epsilon}(2.5+\beta_{i}-2\epsilon)\beta}{1-\alpha_{i}}$. Since $\alpha_{i}$ and $\beta_{i}$ are extrapolation parameters, we first fix them as 00 to find $\tau_{i}$. As analyzed before, $\tau_{i}$ can be decided by $\alpha_{i}$, $\beta_{i}$, and $\beta$. The remaining algorithm parameters are $\tau_{i}$, $i=1,2$, and decided by model parameter $\beta$. Hence, we first study the model parameters $\alpha$ and $\beta$ with fixed $\alpha_{i}$ and $\beta_{i}$. After the optimized $\alpha$ and $\beta$ are obtained, we will focus on the algorithm parameters $\alpha_{i}$ and $\beta_{i}$. We first study $\alpha$ and $\beta$ in interval $[1,200]$ with stepsize $10$. We find that our model is quite robust with parameter $\beta$ while the curve is turbulent alone $\alpha$. Hence we narrow the interval to $[5,50]$ and $[0.1,10]$ use the linespace to find $20$ numbers, respectively. Fig. 7 plots the PSNR and NIQE results along the model parameters $\alpha\in[5,50]$ and $\beta\in[0.1,10]$. The surface indicates that the proposed model is stable to parameters $\alpha$ and $\beta$ in the given interval. Since curves become quite flat as the parameter $\alpha$ from $3$ to $5$ along the $\alpha$ axis, which implies choosing $\alpha$ in the interval $[10,40]$ ensures the best result. Similarly, the interval of model parameter $\beta$ is chosen at $[0.1,5]$.
With the obtained optimized model parameter $\beta$, the algorithm parameter $\tau_{i}$ then controlled by $\alpha_{i}$ and $\beta_{i}$, $i=1,2$. Hence, only $\alpha_{i}\in[0,0.5)$ and $\beta_{i}\in[0,1)$ are studied in the algorithm parameter. For these four algorithm parameters, we choose to tune one parameter in its interval while others are fixed as 00. More specifically, we set the model parameters $\alpha_{i}\in[0.01:0.1:0.5]$ and $\beta_{i}\in[0.01:0.1:1]$. From Fig. 8 and Fig. 9, we know that extrapolation parameters speed up the whole algorithm with convergence.
Figure 7: PSNR ($\uparrow$) and NIQE ($\downarrow$) results along with model parameter $\alpha$ and $\beta$.
Figure 8: Energy with log of algorithm parameters $\alpha_{1}$ and $\alpha_{2}$.
Figure 9: Energy with log of algorithm parameters $\beta_{1}$ and $\beta_{2}$.
V-D Image enhancement
In this section, $15$ low-light images in the dataset RealBlur [49] with paired bright-light images are tested to demonstrate the validity of the proposed approach. Here we indicate these real low-light images as RealBlur15. We compare our method with some state-of-the-art image enhancement models, including the traditional methods (LR3M [12], QSPE [50], DRE [13], RMR [14]) and the deep learning methods (RetinexNet [16], URetinex [18], EFINet [51], RQLLIE [52]).
We present the PSNR and SSIM results of dataset Realblur15 in Table II. The highest and the second highest PSNR and SSIM values are marked as red and blue, respectively. From numerical results, the proposed method with the learned prior is better than the methods with traditional priors in most images. While comparing to the deep learning methods, some results fail to achieve the best one, most of our results still get the second best index. Our method achieves the best averaged overall performance in terms of PSNR and SSIM among all the methods, which shows the effectiveness of our method on real low-light images. Besides the numerical results, the visual results of Img1 from the dataset Realblur15 are presented in Fig 10. The input image is with non-uniformed light and small spark lights, which is the real scene image and is hard to handle. As shown in Fig. 10 (b), (c), and (e), we can see that the small spark lights affect the image enhancement results and make the lights abnormal. Compared to the paired ground-truth image (j), Fig. 10 (b), (d), (f), and (g) have a color bias due to the non-uniformed lights. Although the proposed approach generates the best results among the presented images, the color of (i) is still not exactly the same as the ground-truth image (j). Therefore, image enhancement for real-world images is still a challenging task.
TABLE II: Quantitative results with PSNR and SSIM values on datasets Realblur15. The BEST result is in red whereas the SECOND BEST one is in blue.
Input RetinexNet [16] URetinex [18] EFINet [51] RQLLIE [52] LR3M [12] QSPE [50] DRE [13] RWR [14] Ours
Img1 8.91/0.2466 16.93/0.7078 19.69/0.8665 16.60/0.8372 16.56/0.7948 15.70/0.7926 12.53/0.6727 16.35/0.7780 19.92/0.8498 20.36/0.8723
Img2 8.47/0.2965 14.41/0.6872 25.16/0.8729 15.36/0.7933 13.12/0.6481 14.33/0.7368 11.52/0.6803 15.35/0.7098 19.13/0.8352 18.61/0.8596
Img3 9.49/0.2397 16.82/0.6897 23.76/0.8883 16.34/0.8038 15.36/0.7621 15.98/0.7934 13.22/0.6777 16.75/0.8222 17.85/0.8016 20.33/0.8702
Img4 13.09/0.2390 15.07/0.4498 13.63/0.5800 21.76/0.8239 13.56/0.5715 19.19/0.7236 15.83/0.5378 13.99/0.5171 15.85/0.5985 19.26/0.6801
Img5 10.16/0.2381 15.25/0.7264 17.13/0.8159 16.14/0.8683 17.16/0.8158 18.80/0.8760 13.29/0.6129 14.83/0.8171 14.14/0.7539 19.01/0.8826
Img6 6.71/0.3467 14.01/0.6451 21.61/0.8886 14.82/0.7478 10.72/0.6395 12.20/0.7066 9.22/0.6239 11.71/0.6663 15.05/0.8354 15.98/0.8307
Img7 10.44/0.2075 15.28/0.5945 15.56/0.7664 19.54/0.8835 25.13/0.8608 14.09/0.7214 22.66/0.7505 24.97/0.6755 13.97/0.6918 21.84/0.8677
Img8 8.56/0.2137 17.30/0.7930 18.74/0.8919 16.78/0.8919 26.03/0.9200 14.78/0.8682 19.41/0.5935 23.02/0.8793 11.32/0.5343 22.75/0.9574
Img9 10.38/0.1908 16.05/0.6807 14.82/0.7570 18.56/0.8778 22.70/0.8832 18.23/0.8832 21.62/0.6247 19.20/0.7080 13.91/0.6787 21.51/0.8801
Img10 10.43/0.2186 15.82/0.6982 14.95/0.8126 20.89/0.9346 18.04/0.8316 15.33/0.8019 20.38/0.5057 17.12/0.8373 13.64/0.5881 22.80/0.9131
Img11 10.58/0.2415 14.66/0.5793 14.89/0.7024 18.06/0.8365 21.91/0.8894 15.97/0.8108 16.69/0.7629 18.70/0.7202 17.75/0.8643 19.86/0.8331
Img12 10.90/0.2256 16.69/0.5848 14.17/0.7069 18.64/0.8678 25.52/0.8980 19.39/0.7947 17.51/0.5001 21.27/0.6521 13.74/0.6264 20.54/0.7987
Img13 11.87/0.1596 15.57/0.6384 14.03/0.7377 20.33/0.8734 16.95/0.7662 21.76/0.8336 17.54/0.4619 23.34/0.8198 14.89/0.5067 22.94/0.8586
Img14 9.93/0.2665 14.44/0.5629 18.84/0.7712 17.28/0.8079 17.10/0.7608 16.68/0.7656 14.87/0.6558 17.49/0.7638 20.39/0.8260 18.85/0.7755
Img15 9.30/0.2724 14.86/0.6373 17.67/0.7925 16.29/0.8310 16.05/0.7805 15.99/0.7836 12.36/0.6564 15.42/0.7802 16.25/0.7767 18.55/0.8429
Ave. 9.95/0.2402 15.54/0.6450 17.64/0.7900 17.83/0.8452 18.39/0.7881 16.56/0.7928 15.91/0.6211 17.97/0.7431 15.85/0.71787 20.21/0.8482
(a) Input (10.16/0.2381)
(b) RetinexNet (15.25/0.7264)
(c) EFINet (16.14/0.8683)
(d) RQLLIE (17.16/0.8158)
(e) LR3M (18.80/0.8760)
(f) QSPE (13.29/0.6129)
(g) DRE (14.83/0.8171)
(h) RWR (14.14/0.7539)
(i) Ours (19.01/0.8826)
(j) GT
[…]
TABLE III: Average image enhancement results on non-reference datasets DICM, LIME, and MEF. We highlighted the BEST result in red and the SECOND BEST one in blue.
Datasets Index EFINet Bread RQLLIE LR3M DRE RWR Ours
[…]
(a) Input
(b) EFINet (2.247)
(c) Bread (2.535)
(d) RQLLIE (3.144)
(e) LR3M (3.403)
(f) DRE (2.144)
(g) RWR (3.365)
(h) Ours (2.139)
[…]
(a) Input
(b) EFINet (4.220)
(c) Bread (5.132)
(d) RQLLIE (4.424)
(e) LR3M (4.801)
(f) DRE (4.586)
(g) RWR (5.244)
(h) Ours (3.519)
[…]
(a) Input
(b) EFINet (3.333)
(c) Bread (3.779)
(d) RQLLIE (4.161)
(e) LR3M (3.935)
(f) DRE (3.597)
(g) RWR (4.054)
(h) Ours (3.224)
[…]
V-E Non-reference quality assessment
In this subsection, three classical real-world non-reference datasets, DICM [54], LIME [55], and MEF [56] are tested to verify the robustness of the proposed scheme. The state-of-the-art methods EFINet [51], Bread [53], RQLLIE [52], LR3M [12], DRE [13], RMR [14] are compared. Due to the non-reference of the images, we conduct the non-reference metric ARISM, NIQE, and FADE as our index. Note that the higher the ARISM and the lower the NIQE and FADE, the better the image quality is. In Table. III, the average numerical results of the three datasets are listed. With the help of the learned prior, the proposed method has a better performance compared to the approaches with the traditional priors. For most circumstances, the proposed methods have a competitive result.
In Fig. 11, we present the image enhancement results of the image in dataset DICM. The input image (a) is quite simple to handle compared to other extreme dark images. However, the zoom part of results (c), (d), and (g) remain dark. Compared to other methods, our approach generates the enhanced image (h) with a normal light. An image from dataset LIME with edge feature is presented in Fig. 12. For better understanding, we zoom the texture part, which is a keyboard. Our result (h) has clear details of the keyboard in the enhanced image. From the zoom part of the input image (a), the keyboard is under the non-uniformed light, hence (b), (c), (e), and (f) are with the color bias. For (d) and (g), though the color remains, the details cannot be well preserved. In Fig. 13, we present the image enhancement results of the image in Dataset MEF. Similar color bias happens in (d), (e), and (g). For better visuality, we zoom in on the mug in the image. Due to the help of the proposed edge network, our result (h) has a better enhancement performance while others still have low contrast. In conclusion, the proposed method achieved competitive results in enhancing real-world low-light images, as evidenced by both numerical and visual evaluations. The design of the learned edge prior makes the proposed method particularly effective in handling texture images.
VI Conclusion
This paper proposed an edge-guided Retinex model with a novel inertial Bregman alternating linear algorithm for low-light image enhancement. With the help of deep learning, we trained the edge network with paired low-light images with multi-scale residual blocks to obtain the edge prior. The learned edge prior allows us to directly capture edge features from the low-light image, which helps the image enhancement. Additionally, we designed an efficient inertial Bregman alternating linearized minimization algorithm to tackle the proposed Retinex model. The effectiveness of the proposed algorithm was demonstrated through both theoretical and experimental analysis. To better illustrate the superiority of the proposed scheme, some state-of-the-art models on real-world datasets were compared. Based on the presented experimental results, the proposed method with learned prior and solved by the proposed algorithm achieves competitive performance in enhancing the low-light images. However, we acknowledge that our results still exhibit a color bias issue when dealing with non-uniform real-world low-light images. To address this, we plan to design an efficient model in the future to improve the performance of non-uniformed low-light image enhancement.
Appendix A Proof of Lemma 1
To prove Lemma 1, we first recall the descent property of the smooth function in the following lemma.
Lemma 3.
[35] Let $f:~{}\mathbb{R}^{n}\rightarrow\mathbb{R}$ be a continuously differentiable function with gradient $\nabla f$ assumed $L_{f}$-Lipschitz continuous. Then, for any $u,v\in\mathbb{R}^{n}$, we have
$\left|f(u)-f(v)-\langle u-v,\nabla f(v)\rangle\right|\leq\frac{L_{f}}{2}\|u-v\|^{2}.$
In the following, we prove Lemma 1.
Proof.
Using the updating schemes of $x_{1}^{k+1}$ and $x_{2}^{k+1}$ in (15c) and (16c), respectively, we have
$\theta_{1}\left(x_{1}^{k+1}\right)+h^{+}\left(x_{1}^{k+1},x_{2}^{k}\right)+\tau_{1}^{k}D_{\psi_{k}}\left(x_{1}^{k+1},x_{1}^{k}\right)$
$-\left\langle x_{1}^{k+1}\right.\left.-x_{1}^{k},\nabla_{x_{1}}h^{-}\left(z_{1}^{k},x_{2}^{k}\right)+\tau_{1}^{k}\left(y_{1}^{k}-x_{1}^{k}\right)\right\rangle$
$\leq\theta_{1}\left(x_{1}^{k}\right)+h^{+}\left(x_{1}^{k},x_{2}^{k}\right),$
and
$\theta_{2}\left(x_{2}^{k+1}\right)+h^{+}\left(x_{1}^{k+1},x_{2}^{k+1}\right)+\tau_{2}^{k}D_{{\varphi_{k}}}\left(x_{2}^{k+1},x_{2}^{k}\right)$
$-\left\langle x_{2}^{k+1}-x_{2}^{k},\nabla_{x_{2}}h^{-}\left(x_{1}^{k+1},z_{2}^{k}\right)+\tau_{2}^{k}\left(y_{2}^{k}-x_{2}^{k}\right)\right\rangle$
$\leq\theta_{2}\left(x_{2}^{k}\right)+h^{+}\left(x_{1}^{k+1},x_{2}^{k}\right).$
Summing both inequalities and using the definition of $\Phi(\cdot,\cdot)$ in (13) yield
$\Phi\left(x_{1}^{k+1},x_{2}^{k+1}\right)$
$\leq$ $\Phi\left(x_{1}^{k},x_{2}^{k}\right)-h^{-}\left(x_{1}^{k+1},x_{2}^{k+1}\right)+h^{-}\left(x_{1}^{k},x_{2}^{k}\right)$
$+\left\langle x_{1}^{k+1}\right.\left.-x_{1}^{k},\nabla_{x_{1}}h^{-}\left(z_{1}^{k},x_{2}^{k}\right)+\tau_{1}^{k}\left(y_{1}^{k}-x_{1}^{k}\right)\right\rangle$
$+\left\langle x_{2}^{k+1}-x_{2}^{k},\nabla_{x_{2}}h^{-}\left(x_{1}^{k+1},z_{2}^{k}\right)+\tau_{2}^{k}\left(y_{2}^{k}-x_{2}^{k}\right)\right\rangle$
$-\tau_{1}^{k}D_{\psi_{k}}\left(x_{1}^{k+1},x_{1}^{k}\right)-\tau_{2}^{k}D_{{\varphi_{k}}}\left(x_{2}^{k+1},x_{2}^{k}\right).$
From Assumption 2(i) and Lemma 3, we know that
$-h^{-}\left(x_{1}^{k+1},x_{2}^{k}\right)\leq$ $\frac{L_{1}^{-}\left(x_{2}^{k}\right)}{2}\left\|x_{1}^{k+1}-x_{1}^{k}\right\|^{2}-h^{-}\left(x_{1}^{k},x_{2}^{k}\right)$
$-\left\langle x_{1}^{k+1}-x_{1}^{k},\nabla_{x_{1}}h^{-}\left(x_{1}^{k},x_{2}^{k}\right)\right\rangle,$
where $L_{1}^{-}\left(x_{2}^{k}\right)$ is the Lipschitz constant of $\nabla_{x_{1}}h^{-}\left(x_{1},x_{2}^{k}\right)$ at $x_{1}^{k}$. Similarly, we have
$-h^{-}\left(x_{1}^{k+1},x_{2}^{k+1}\right)$
$\leq$ $\frac{L_{2}^{-}\left(x_{1}^{k+1}\right)}{2}\left\|x_{2}^{k+1}-x_{2}^{k}\right\|^{2}-h^{-}\left(x_{1}^{k+1},x_{2}^{k}\right)$
$-\left\langle x_{2}^{k+1}-x_{2}^{k},\nabla_{x_{2}}h^{-}\left(x_{1}^{k+1},x_{2}^{k}\right)\right\rangle,$
where $L_{2}^{-}\left(x_{1}^{k+1}\right)$ is the Lipschitz constant of $\nabla_{x_{2}}h^{-}\left(x_{1}^{k+1},x_{2}\right)$ at $x_{2}^{k}$. Summing the above two inequalities and plugging into (A), we get
$\Phi\left(x_{1}^{k+1},x_{2}^{k+1}\right)$ (27)
$\leq$ $\Phi\left(x_{1}^{k},x_{2}^{k}\right)-\tau_{1}^{k}D_{\psi_{k}}\left(x_{1}^{k+1},x_{1}^{k}\right)-\tau_{2}^{k}D_{{\varphi_{k}}}\left(x_{2}^{k+1},x_{2}^{k}\right)$
$+\left\langle x_{1}^{k+1}-x_{1}^{k},\nabla_{x_{1}}h^{-}\left(z_{1}^{k},x_{2}^{k}\right)-\nabla_{x_{1}}h^{-}\left(x_{1}^{k},x_{2}^{k}\right)\right.$
$+\left\langle x_{2}^{k+1}-x_{2}^{k},\nabla_{x_{2}}h^{-}\left(x_{1}^{k+1},z_{2}^{k}\right)-\nabla_{x_{2}}h^{-}\left(x_{1}^{k+1},x_{2}^{k}\right)\right.$
$+\tau_{1}^{k}\left(y_{1}^{k}-x_{1}^{k}\right)+\tau_{2}^{k}\left(y_{2}^{k}-x_{2}^{k}\right)$
$+\frac{L_{1}^{-}\left(x_{2}^{k}\right)}{2}\left\|x_{1}^{k+1}-x_{1}^{k}\right\|^{2}+\frac{L_{2}^{-}\left(x_{1}^{k+1}\right)}{2}\left\|x_{2}^{k+1}-x_{2}^{k}\right\|^{2}.$
Note that
$\left\langle x_{1}^{k+1}\right.\left.-x_{1}^{k},\nabla_{x_{1}}h^{-}\left(z_{1}^{k},x_{2}^{k}\right)-\nabla_{x_{1}}h^{-}\left(x_{1}^{k},x_{2}^{k}\right)\right\rangle$ (28)
$\leq\frac{L_{1}^{-}(x_{2}^{k})\beta_{1}^{k}}{2}\left\|x_{1}^{k+1}-x_{1}^{k}\right\|^{2}$
$\quad+\frac{1}{2L_{1}^{-}(x_{2}^{k})\beta_{1}^{k}}\left\|\nabla_{x_{1}}h^{-}\left(z_{1}^{k},x_{2}^{k}\right)-\nabla_{x_{1}}h^{-}\left(x_{1}^{k},x_{2}^{k}\right)\right\|$
$\leq\frac{L_{1}^{-}(x_{2}^{k})\beta_{1}^{k}}{2}\left\|x_{1}^{k+1}-x_{1}^{k}\right\|^{2}+\frac{L_{1}^{-}(x_{2}^{k})}{2\beta_{1}^{k}}\left\|z_{1}^{k}-x_{1}^{k}\right\|^{2}.$
This together with (15a) and (15b), we have
$\left\langle x_{1}^{k+1}\right.\left.-x_{1}^{k},\nabla_{x_{1}}h^{-}\left(z_{1}^{k},x_{2}^{k}\right)-\nabla_{x_{1}}h^{-}\left(x_{1}^{k},x_{2}^{k}\right)+\tau_{1}^{k}\left(y_{1}^{k}-x_{1}^{k}\right)\right\rangle$ (29)
$\leq\left(\frac{L_{1}^{-}(x_{2}^{k})\beta_{1}^{k}}{2}+\frac{\alpha_{1}^{k}\tau_{1}^{k}}{2}\right)\left\|x_{1}^{k+1}-x_{1}^{k}\right\|^{2}$
$\quad+\left(\frac{L_{1}^{-}(x_{2}^{k})\beta_{1}^{k}}{2}+\frac{\alpha_{1}^{k}\tau_{1}^{k}}{2}\right)\left\|x_{1}^{k}-x_{1}^{k-1}\right\|^{2}.$
Similarly,
$\left\langle x_{2}^{k+1}-x_{2}^{k},\nabla_{x_{2}}h^{-}\left(x_{1}^{k+1},z_{2}^{k}\right)-\nabla_{x_{2}}h^{-}\left(x_{1}^{k+1},x_{2}^{k}\right)+\tau_{2}^{k}\left(y_{2}^{k}-x_{2}^{k}\right)\right\rangle$ (30)
$\leq\left(\frac{L_{2}^{-}(x_{1}^{k+1})\beta_{2}^{k}}{2}+\frac{\alpha_{2}^{k}\tau_{2}^{k}}{2}\right)\left\|x_{2}^{k+1}-x_{2}^{k}\right\|^{2}$
$\quad+\left(\frac{L_{2}^{-}(x_{1}^{k+1})\beta_{2}^{k}}{2}+\frac{\alpha_{2}^{k}\tau_{2}^{k}}{2}\right)\left\|x_{2}^{k}-x_{2}^{k-1}\right\|^{2}.$
Recalling the strong convexity of ${\psi_{k}}(\cdot)$ and ${\varphi_{k}}(\cdot)$, we accordingly chose ${\psi_{k}}(\cdot)$ and ${\varphi_{k}}(\cdot)$ with strongly convex modulus $\rho_{1}$ and $\rho_{2}$, respectively, so that
$D_{{\psi_{k}}}\left(x_{1}^{k+1},x_{1}^{k}\right)\geq\frac{\rho_{1}}{2}\left\|x_{1}^{k+1}-x_{1}^{k}\right\|^{2},$ (31)
$D_{\varphi_{k}}\left(x_{2}^{k+1},x_{2}^{k}\right)\geq\frac{\rho_{2}}{2}\left\|x_{2}^{k+1}-x_{2}^{k}\right\|^{2}.$
Substituting (29), (30) and (31) into (27), we obtain
$\Phi\left(x_{1}^{k+1},x_{2}^{k+1}\right)-\Phi\left(x_{1}^{k},x_{2}^{k}\right)$ (32)
$\leq-\left(\frac{\rho_{1}\tau_{1}^{k}}{2}-\frac{L_{1}^{-}\left(x_{2}^{k}\right)}{2}-\frac{L_{1}^{-}(x_{2}^{k})\beta_{1}^{k}}{2}-\frac{\alpha_{1}^{k}\tau_{1}^{k}}{2}\right)\left\|x_{1}^{k+1}-x_{1}^{k}\right\|^{2}$
$\quad+\left(\frac{L_{1}^{-}(x_{2}^{k})\beta_{1}^{k}}{2}+\frac{\alpha_{1}^{k}\tau_{1}^{k}}{2}\right)\left\|x_{1}^{k}-x_{1}^{k-1}\right\|^{2}$
$\quad-\left(\frac{\rho_{2}\tau_{2}^{k}}{2}-\frac{L_{2}^{-}\left(x_{1}^{k+1}\right)}{2}-\frac{L_{2}^{-}(x_{1}^{k+1})\beta_{2}^{k}}{2}-\frac{\alpha_{2}^{k}\tau_{2}^{k}}{2}\right)\left\|x_{2}^{k+1}-x_{2}^{k}\right\|^{2}$
$\quad+\left(\frac{L_{2}^{-}(x_{1}^{k+1})\beta_{2}^{k}}{2}+\frac{\alpha_{2}^{k}\tau_{2}^{k}}{2}\right)\left\|x_{2}^{k}-x_{2}^{k-1}\right\|^{2}.$
According to Assumptions 2 and 3 and the definitions of $\tau_{1}^{k}$ and $\delta_{1}$ in (21) and (22), respectively, we have
$\rho_{1}\tau_{1}^{k}-L_{1}^{-}\left(x_{2}^{k}\right)-L_{1}^{-}(x_{2}^{k})\beta_{1}^{k}-\alpha_{1}^{k}\tau_{1}^{k}-\delta_{1}=(1+\varepsilon)\delta_{1}-\delta_{1}=\varepsilon\delta_{1},$
and
$\delta_{1}-L_{1}^{-}(x_{2}^{k})\beta_{1}^{k}-\alpha_{1}^{k}\tau_{1}^{k}$
$=$ $\ \delta_{1}-\frac{(1+\varepsilon)\delta_{1}+(1+\beta_{1}^{k})L_{1}^{-}(x_{2}^{k})}{\rho_{1}-\alpha_{1}^{k}}\alpha_{1}^{k}-L_{1}^{-}(x_{2}^{k})\beta_{1}^{k}$
$\geq$ $\ \delta_{1}-(1-\varepsilon)\delta_{1}=\varepsilon\delta_{1}.$
Similarly, for the $\tau_{2}^{k}$ and $\delta_{2}$ defined in (21) and (22), it holds that
$\rho_{2}\tau_{2}^{k}-L_{2}^{-}\left(x_{1}^{k+1}\right)-L_{2}^{-}(x_{1}^{k+1})\beta_{2}^{k}-\alpha_{2}^{k}\tau_{2}^{k}-\delta_{2}=\varepsilon\delta_{2},$
and
$\delta_{2}-L_{2}^{-}(x_{1}^{k+1})\beta_{2}^{k}-\alpha_{2}^{k}\tau_{2}^{k}=\delta_{2}-(1-\varepsilon)\delta_{2}^{k}\geq\varepsilon\delta_{2}.$
Plugging them into (32), we have
$\Phi\left(x_{1}^{k+1},x_{2}^{k+1}\right)+\frac{\delta_{1}}{2}\|x_{1}^{k+1}-x_{1}^{k}\|^{2}+\frac{\delta_{2}}{2}\|x_{2}^{k+1}-x_{2}^{k}\|^{2}$ (33)
$\leq\Phi\left(x_{1}^{k},x_{2}^{k}\right)+\frac{\delta_{1}}{2}\|x_{1}^{k}-x_{1}^{k-1}\|^{2}+\frac{\delta_{2}}{2}\|x_{2}^{k}-x_{2}^{k-1}\|^{2}$
$\leq-\frac{\varepsilon\delta_{1}}{2}\left(\left\|x_{1}^{k+1}-x_{1}^{k}\right\|^{2}+\left\|x_{1}^{k}-x_{1}^{k-1}\right\|^{2}\right)$
$\quad-\frac{\varepsilon\delta_{2}}{2}\left(\left\|x_{2}^{k+1}-x_{2}^{k}\right\|^{2}+\left\|x_{2}^{k}-x_{2}^{k-1}\right\|^{2}\right).$
This together the definition of $\Theta_{\delta_{1},\delta_{2}}$ in (24) and $\delta=\frac{\varepsilon}{2}\min\{\delta_{1},\delta_{2}\}$, we can obtain (25) easily. This completes the proof. ∎
Appendix B Proof of Lemma 2
Proof.
According to (14), the optimality condition of the subproblem (15c) can be read as
$0\in$ $\partial\theta_{1}(x_{1}^{k+1})+\nabla_{x_{1}}h^{+}(x_{1}^{k+1},x_{2}^{k})-\nabla_{x_{1}}h^{-}(z_{1}^{k},x_{2}^{k})$ (34)
$-\tau_{1}^{k}(y_{1}^{k}-x_{1}^{k})+\tau_{1}^{k}(\nabla{\psi_{k}}(x_{1}^{k+1})-\nabla{\psi_{k}}(x_{1}^{k})).$
Similarly, the optimality condition of the subproblem (16c) can be written as
$0\in$ $\partial\theta_{2}(x_{2}^{k+1})+\nabla_{x_{2}}h^{+}(x_{1}^{k+1},x_{2}^{k+1})-\nabla_{x_{2}}h^{-}(x_{1}^{k+1},z_{2}^{k})$ (35)
$-\tau_{2}^{k}(y_{2}^{k}-x_{2}^{k})+\tau_{2}^{k}(\nabla\varphi(x_{2}^{k+1})-\nabla\varphi(x_{2}^{k})).$
First, from the definition of $\Theta_{\delta_{1},\delta_{2}}$ in (24), we have
$\partial_{w_{11}}\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k+1})$ (36)
$=\partial\theta_{1}(x_{1}^{k+1})+\nabla_{x_{1}}h^{+}(x_{1}^{k+1},x_{2}^{k+1})-\nabla_{x_{1}}h^{-}(x_{1}^{k+1},x_{2}^{k+1})$
$\quad+\delta_{1}(x_{1}^{k+1}-x_{1}^{k})$
$\ni\nabla_{x_{1}}h^{+}(x_{1}^{k+1},x_{2}^{k+1})-\nabla_{x_{1}}h^{+}(x_{1}^{k+1},x_{2}^{k})$
$\quad+\nabla_{x_{1}}h^{-}(z_{1}^{k},x_{2}^{k})-\nabla_{x_{1}}h^{-}(x_{1}^{k+1},x_{2}^{k+1})$
$\quad+\tau_{1}^{k}(y_{1}^{k}-x_{1}^{k})+\delta_{1}(x_{1}^{k+1}-x_{1}^{k})$
$\quad-\tau_{1}^{k}(\nabla{\psi_{k}}(x_{1}^{k+1})-\nabla{\psi_{k}}(x_{1}^{k})),$
where the last inclusion follows from (34). Second, we compute the subgradient of $\Theta_{\delta_{1},\delta_{2}}$ with respect to $w_{12}$,
$\partial_{w_{12}}\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k+1})$ (37)
$=\partial\theta_{2}(x_{2}^{k+1})+\nabla_{x_{2}}h^{+}(x_{1}^{k+1},x_{2}^{k+1})-\nabla_{x_{2}}h^{-}(x_{1}^{k+1},x_{2}^{k+1})$
$\quad+\delta_{2}(x_{2}^{k+1}-x_{2}^{k})$
$\ni\nabla_{x_{2}}h^{-}(x_{1}^{k+1},z_{2}^{k})-\nabla_{x_{2}}h^{-}(x_{1}^{k+1},x_{2}^{k+1})$
$\quad+\tau_{2}^{k}(y_{2}^{k}-x_{2}^{k})+\delta_{2}(x_{2}^{k+1}-x_{2}^{k})-\tau_{2}^{k}(\nabla\varphi(x_{1}^{k+1})-\nabla\varphi(x_{1}^{k})),$
where the inclusion follows from (35). Finally, it follows from (24) that
$\nabla_{w_{21}}\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k+1})=\delta_{1}\left(x_{1}^{k}-x_{1}^{k+1}\right),$ (38)
$\nabla_{w_{22}}\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k+1})=\delta_{2}\left(x_{2}^{k}-x_{2}^{k+1}\right).$
Since $\nabla h$ is Lipschitz continuous on bounded subsets of $\mathbb{R}^{n_{1}}\times\mathbb{R}^{n_{2}}$ from Assumption 2, it follows from (36) that
${\rm dist}(0,\partial_{w_{11}}\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k+1}))$
$\leq$ $\|\nabla_{x_{1}}h^{+}(x_{1}^{k+1},x_{2}^{k+1})-\nabla_{x_{1}}h^{+}(x_{1}^{k+1},x_{2}^{k})\|$
$+\|\nabla_{x_{1}}h^{-}(z_{1}^{k},x_{2}^{k})-\nabla_{x_{1}}h^{-}(x_{1}^{k+1},x_{2}^{k+1})\|$
$+\tau_{1}^{k}\|y_{1}^{k}-x_{1}^{k}\|+\delta_{1}\|x_{1}^{k+1}-x_{1}^{k}\|$
$+\tau_{1}^{k}\|\nabla{\psi_{k}}(x_{1}^{k+1})-\nabla{\psi_{k}}(x_{1}^{k})\|$
$\leq$ $L^{+}\|x_{2}^{k+1}-x_{2}^{k}\|+L^{-}\|z_{1}^{k}-x_{1}^{k+1}\|+L^{-}\|x_{2}^{k}-x_{2}^{k+1}\|$
$+\tau_{1}^{k}\alpha_{1}^{k}\|x_{1}^{k}-x_{1}^{k-1}\|+\delta_{1}\|x_{1}^{k+1}-x_{1}^{k}\|+\tau_{1}^{k}L_{{\psi_{k}}}\|x_{1}^{k+1}-x_{1}^{k}\|$
$\leq$ $L\|x_{2}^{k+1}-x_{2}^{k}\|+(\tau_{1}^{+}\bar{\alpha}_{1}+\bar{\beta}_{1}L^{-})\|x_{1}^{k}-x_{1}^{k-1}\|$
$+(\delta_{1}+\tau_{1}^{+}L_{\psi})\|x_{1}^{k+1}-x_{1}^{k}\|,$
where $L=L^{+}+L^{-}$, $L_{\psi}=\sup\{L_{\psi_{k}}:k\in\mathbb{N}\}$, the first and second inequalities follow from Assumption 2, (15a) and (15b), the fact the sequence $\left\{\tau_{1}^{k}\right\}_{k\in\mathbb{N}}$ is bounded from above by $\tau_{1}^{+}$, and $\alpha_{1}^{k}\leq\bar{\alpha}_{1},\beta_{1}^{k}\leq\bar{\beta}_{1}$ for all $k\in\mathbb{N}$. Moreover, it follows from (37) that
${\rm dist}(0,\partial_{w_{12}}\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k+1}))$
$\leq$ $\|\nabla_{x_{2}}h^{-}(x_{1}^{k+1},z_{2}^{k})-\nabla_{x_{2}}h^{-}(x_{1}^{k+1},x_{2}^{k+1})\|+\tau_{2}^{k}\|y_{2}^{k}-x_{2}^{k}\|$
$+\delta_{2}\|x_{2}^{k+1}-x_{2}^{k}\|+\tau_{2}^{k}\|\nabla\varphi_{k}(x_{2}^{k+1})-\nabla\varphi_{k}(x_{2}^{k})\|$
$\leq$ $\lambda_{2}^{+}\|z_{2}^{k}-x_{2}^{k+1}\|+\tau_{2}^{k}\alpha_{2}^{k}\|x_{2}^{k}-x_{2}^{k-1}\|$
$+\delta_{2}\|x_{2}^{k+1}-x_{2}^{k}\|+\tau_{2}^{k}L_{\varphi_{k}}\|x_{2}^{k+1}-x_{2}^{k}\|$
$\leq$ $\lambda_{2}^{+}\|x_{2}^{k}-x_{2}^{k+1}\|+(\lambda_{2}^{+}\bar{\beta}_{2}+\tau_{2}^{+}\bar{\alpha}_{2})\|x_{2}^{k}-x_{2}^{k-1}\|$
$+(\delta_{2}+\tau_{2}^{+}L_{\varphi})\|x_{2}^{k+1}-x_{2}^{k}\|,$
where $L_{\varphi}=\sup\{L_{\varphi_{k}}:k\in\mathbb{N}\}$. Summing up these estimations, we get from (24) that
${\rm dist}\left(0,\partial\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k+1})\right)$
$\leq{\rm dist}(0,\partial_{w_{11}}\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k+1}))+{\rm dist}(0,\partial_{w_{12}}\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k+1}))$
$\quad+\|\nabla_{w_{21}}\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k+1})\|+\|\nabla_{w_{22}}\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k+1})\|$
$\leq L\|x_{2}^{k+1}-x_{2}^{k}\|+(\tau_{1}^{+}\bar{\alpha}_{1}+\bar{\beta}_{1}L^{-})\|x_{1}^{k}-x_{1}^{k-1}\|$
$\quad+(\delta_{1}+\tau_{1}^{+}L_{{\psi_{k}}})\|x_{1}^{k+1}-x_{1}^{k}\|$
$\quad\lambda_{2}^{+}\|x_{2}^{k}-x_{2}^{k+1}\|+(\lambda_{2}^{+}\bar{\beta}_{2}+\tau_{2}^{+}\bar{\alpha}_{2})\|x_{2}^{k}-x_{2}^{k-1}\|$
$\quad+(\delta_{2}+\tau_{2}^{+}L_{\varphi})\|x_{2}^{k+1}-x_{2}^{k}\|$
$\quad+\delta_{1}\|x_{1}^{k}-x_{1}^{k+1}\|+\delta_{2}\|x_{2}^{k}-x_{2}^{k+1}\|.$
Let $\tilde{\gamma}=\max\{L+\lambda_{2}^{+}+2\delta_{2}+\tau_{2}^{+}L_{\varphi},\tau_{1}^{+}\bar{\alpha}_{1}+\bar{\beta}_{1}L^{-},2\delta_{1}+\tau_{1}^{+}L_{\psi},\lambda_{2}^{+}\bar{\beta}_{2}+\tau_{2}^{+}\bar{\alpha}_{2}\}$, we have
${\rm dist}\left(0,\partial\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k+1})\right)$ (39)
$\leq\tilde{\gamma}\left(\|x_{1}^{k+1}-x_{1}^{k}\|+\|x_{1}^{k}-x_{1}^{k-1}\|\right)$
$\quad+\tilde{\gamma}\left(\|x_{2}^{k+1}-x_{2}^{k}\|+\|x_{2}^{k}-x_{2}^{k-1}\|\right)$
$\leq 2\sqrt{2}\tilde{\gamma}\|{\bf w}^{k+1}-{\bf w}^{k}\|,$
This completes the proof with $\gamma=2\sqrt{2}\tilde{\gamma}$. ∎
Appendix C Proof of Theorem 1
Proof.
We first prove (i). It follow from Lemma 1 that for any $N\in\mathbb{N}$,
$\sum_{k=0}^{N}\delta\left\|\mathbf{w}^{k+1}-\mathbf{w}^{k}\right\|^{2}\leq\Theta_{\delta_{1},\delta_{2}}\left(\mathbf{w}^{0}\right)-\Theta_{\delta_{1},\delta_{2}}\left(\mathbf{w}^{N+1}\right).$ (40)
According to Assumption 1, we know that $\Phi$ is bounded from below, and hence $\Theta_{\delta_{1},\delta_{2}}$ is also bounded from below because $\Theta_{\delta_{1},\delta_{2}}(\cdot,\cdot)\geq\Phi(\cdot)$. Thus, letting $N\rightarrow\infty$ in (40) yields that
$\sum_{k=0}^{\infty}\delta\left\|\mathbf{w}^{k+1}-\mathbf{w}^{k}\right\|^{2}<\infty,$
which means that
$\lim_{k\rightarrow\infty}\left\|\mathbf{w}^{k+1}-\mathbf{w}^{k}\right\|=0.$
This fact together with (25) implies that $\left\|\mathrm{w}^{k}\right\|\rightarrow 0$ as $k\rightarrow\infty$.
Now we prove (ii). Since $\left\{\mathbf{w}^{k}\right\}_{k\in\mathbb{N}}$ is assumed to be bounded, the set $z\left(\mathbf{w}^{0}\right)$ is nonempty. Thus the limit point $\mathbf{w}^{*}=\left(x_{1}^{*},x_{2}^{*},\hat{x}_{1},\hat{x}_{2}\right)$ of the subsequence of $\left\{\mathbf{w}^{k}\right\}_{k\in\mathbb{N}^{*}}$ of $\left\{\mathbf{w}^{k_{j}}\right\}_{j\in\mathbb{N}}$ is exits. We will prove that $\mathbf{w}^{*}$ is a critical point of $\Theta_{\delta_{1},\delta_{2}}$ defined in (24). It follows from (15c) and $k=k_{j}-1$ that
$\theta_{1}\left(x_{1}^{k_{j}}\right)+h^{+}(x_{1}^{k_{j}},x_{2}^{k_{j}-1})+\tau_{1}^{k_{j}-1}D_{\psi}(x_{1}^{k_{j}},x_{1}^{k_{j}-1})$ (41)
$\leq\theta_{1}\left(x_{1}^{*}\right)+h^{+}(x_{1}^{*},x_{2}^{k_{j}-1})+\tau_{1}^{k_{j}-1}D_{\psi}(x_{1}^{*},x_{1}^{k_{j}-1})$
$\quad-\left\langle x_{1}^{*}-x_{1}^{k_{j}},\nabla_{x_{1}}h^{-}(z_{1}^{k_{j}-1},x_{2}^{k_{j}-1})+\tau_{1}^{k_{j}-1}(y_{1}^{k_{j}-1}-x_{1}^{k_{j}-1})\right\rangle.$
Similarly, it follows from (16c) that
$\theta_{2}\left(x_{2}^{k_{j}}\right)+h^{+}(x_{1}^{k_{j}},x_{2}^{k_{j}})+\tau_{2}^{k_{j}-1}D_{\phi}(x_{2}^{k_{j}},x_{2}^{k_{j}-1})$ (42)
$\leq\theta_{2}\left(x_{2}^{*}\right)+h^{+}(x_{1}^{k_{j}},x_{2}^{*})+\tau_{2}^{k_{j}-1}D_{\phi}(x_{2}^{*},x_{2}^{k_{j}-1})$
$\quad-\left\langle x_{2}^{*}-x_{2}^{k_{j}},\nabla_{x_{2}}h^{-}(x_{1}^{k_{j}},z_{2}^{k_{j}-1})+\tau_{2}^{k_{j}-1}(y_{2}^{k_{j}-1}-x_{2}^{k_{j}-1})\right\rangle.$
Hence, it follows from the nonnegativity of Bregman distance and the inequalities (41) and (42) that
$\lim_{j\rightarrow\infty}\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k})$ (43)
$=\lim_{j\rightarrow\infty}\Big{\{}\theta_{1}\left(x_{1}^{k_{j}}\right)+\theta_{2}\left(x_{2}^{k_{j}}\right)+h\left(x_{1}^{k_{j}},x_{2}^{k_{j}}\right)$
$\quad+\frac{\delta_{1}}{2}\left\|x_{1}^{k_{j}}-x_{1}^{k_{j}-1}\right\|^{2}+\frac{\delta_{2}}{2}\left\|x_{2}^{k_{j}}-x_{2}^{k_{j}-1}\right\|^{2}\Big{\}}$
$\leq\lim_{j\rightarrow\infty}\Big{\{}\theta_{1}\left(x_{1}^{k_{j}}\right)+\theta_{2}\left(x_{2}^{k_{j}}\right)+h\left(x_{1}^{k_{j}},x_{2}^{k_{j}}\right)$
$\quad+\frac{\delta_{1}}{2}\left\|x_{1}^{k_{j}}-x_{1}^{k_{j}-1}\right\|^{2}+\frac{\delta_{2}}{2}\left\|x_{2}^{k_{j}}-x_{2}^{k_{j}-1}\right\|^{2}$
$\quad+\tau_{1}^{k_{j}-1}D_{\psi}(x_{1}^{k_{j}},x_{1}^{k_{j}-1})+\tau_{2}^{k_{j}-1}D_{\psi}(x_{2}^{k_{j}},x_{2}^{k_{j}-1})\Big{\}}$
$\leq\limsup_{j\rightarrow\infty}\Big{\{}\theta_{1}\left(x_{1}^{*}\right)+h^{+}(x_{1}^{*},x_{2}^{k_{j}-1})$
$\quad+\tau_{1}^{k_{j}-1}D_{\psi}(x_{1}^{*},x_{1}^{k_{j}-1})-h^{+}(x_{1}^{k_{j}},x_{2}^{k_{j}-1})$
$\quad-\left\langle x_{1}^{*}-x_{1}^{k_{j}},\nabla_{x_{1}}h^{-}(z_{1}^{k_{j}-1},x_{2}^{k_{j}-1})+\tau_{1}^{k_{j}-1}(y_{1}^{k_{j}-1}-x_{1}^{k_{j}-1})\right\rangle$
$\quad+\frac{\delta_{1}}{2}\left\|x_{1}^{k_{j}}-x_{1}^{k_{j}-1}\right\|^{2}+\theta_{2}\left(x_{2}^{*}\right)$
$\quad+h^{+}(x_{1}^{k_{j}},x_{2}^{*})+\tau_{2}^{k_{j}-1}D_{\phi}(x_{2}^{*},x_{2}^{k_{j}-1})-h^{-}(x_{1}^{k_{j}},x_{2}^{k_{j}})$
$\qquad-\left\langle x_{2}^{*}-x_{2}^{k_{j}},\nabla_{x_{2}}h^{-}(x_{1}^{k_{j}},z_{2}^{k_{j}-1})+\tau_{2}^{k_{j}-1}(y_{2}^{k_{j}-1}-x_{2}^{k_{j}-1})\right\rangle$
$\quad+\frac{\delta_{2}}{2}\left\|x_{2}^{k_{j}}-x_{2}^{k_{j}-1}\right\|^{2}\Big{\}}$
$=\theta_{1}(x_{1}^{*})+h(x_{1}^{*},x_{2}^{*})+\theta_{2}(x_{2}^{*})\leq\Theta_{\delta_{1},\delta_{2}}({\bf w}^{*})=\Theta^{*}.$
where the second equality follows from the continuity of $h(\cdot,\cdot)$ and $\lim_{j\rightarrow\infty}\left\|\boldsymbol{w}^{k_{j}}-\boldsymbol{w}^{k_{j}-1}\right\|=0$ proved in (i). From Assumption 1, we have
$\liminf_{j\rightarrow\infty}\theta_{1}\left(x_{1}^{k_{j}}\right)\geq\theta_{1}\left(x_{1}^{*}\right)\text{ and }\liminf_{j\rightarrow\infty}\theta_{2}\left(x_{2}^{k_{j}}\right)\geq\theta_{2}\left(x_{2}^{*}\right).$
The above relations yield
$\Theta^{*}\equiv\Theta_{\delta_{1},\delta_{2}}\left(\bf{w}^{*}\right)\leq\liminf_{j\rightarrow\infty}\Theta_{\delta_{1},\delta_{2}}\left({\bf w}^{k_{j}}\right).$
Consequently, $\lim_{j\rightarrow\infty}\Theta_{\delta_{1},\delta_{2}}\left({\bf w}^{k_{j}}\right)=\Theta^{*}$. On the other hand, it follows from Lemmas 2 that
$\lim_{j\rightarrow\infty}\operatorname{dist}\left(0,\partial\Theta_{\delta_{1},\delta_{2}}\left(\bf{w}^{k_{j}}\right)\right)\leq\lim_{j\rightarrow\infty}\gamma\left\|\bf{w}^{k_{j}}-\bf{w}^{k_{j}-1}\right\|=0.$
Now, the closedness property of $\partial\Theta_{\delta_{1},\delta_{2}}$ refers that $0\in\partial\Psi\left(\mathbf{w}^{*}\right)$, which proves that $\mathbf{w}^{*}$ is a critical point of $\Theta_{\delta_{1},\delta_{2}}$.
Now we prove (iii). According to Assumption 1, the sequence $\{\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k})\}$ is bounded below. Besides, the sequence $\{\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k})\}$ is nonincreasing from (25). Thus, the limit $\lim_{k\rightarrow\infty}\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k})=\Theta^{\infty}$ exists. Taking any ${\bf w}^{*}\in z({\bf w}^{0})$ as an accumulation point of $\{{\bf w}^{k}\}$, then there exists a subsequence $\{{\bf w}^{k_{j}}\}$ satisfying $\lim_{j\rightarrow\infty}{\bf w}^{k_{j}}={\bf w}^{*}$. From Items (i) and (ii), we have
$\Theta_{\delta_{1},\delta_{2}}({\bf w}^{*})=\lim_{j\rightarrow\infty}\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k_{j}})=\lim_{k\rightarrow\infty}\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k})=\Theta^{\infty}.$
Thus, we complete the proof by the arbitrariness of ${\bf w}^{*}$ in $z({\bf w}^{0})$. ∎
Appendix D Proof of Theorem 2
To prove Theorem 2, we need to introduce a uniformized KL property which was established in [35] in the following, it will be useful for further convergence analysis.
Lemma 4.
[35] (Uniformized KL property) Let $\Omega$ be a compact set and $f:\mathbb{R}^{n}\rightarrow(-\infty,+\infty]$ be a proper and lower semicontinuous function. Assume that $f$ is a constant on $\Omega$ and satisfies the KL property at each point of $\Omega$. Then, there exist $\varsigma>0,~{}\eta>0$ and $\varphi\in\Phi_{\eta}$ such that
$\varphi^{\prime}(f(x)-f(\bar{x})){\rm dist}(0,\partial f(x))\geq 1,$ (44)
for all $\bar{x}\in\Omega$ and each $x$ satisfying ${\rm dist}(x,\Omega)<\varsigma$ and $f(\bar{x})<f(x)<f(\bar{x})+\eta.$
Now we begin to prove Theorem 2.
Proof.
It follows from Theorem 1 that for any ${\bf w}^{*}\in z({\bf w}^{0})$, there exists a subsequence $\{{\bf w}^{k_{j}}\}$ of $\{{\bf w}^{k}\}$ converging to ${\bf w}^{*}$, and $\lim_{k\rightarrow\infty}\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k})=\Theta_{\delta_{1},\delta_{2}}({\bf w}^{*})$. If there exists an integer $\bar{k}$ such that $\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k})=\Theta_{\delta_{1},\delta_{2}}({\bf w}^{*})$, then from Lemma 1, we have
$\delta\left\|\mathbf{w}^{k+1}-\mathbf{w}^{k}\right\|^{2}$ $\leq\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k})-\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k+1})$
$\leq\Theta_{\delta_{1},\delta_{2}}({\bf w}^{\bar{k}})-\Theta_{\delta_{1},\delta_{2}}({\bf w}^{*})=0,~{}~{}\forall k>\bar{k}.$
Thus, we have ${\bf w}^{k+1}={\bf w}^{k}$ for any $k>\bar{k}$. Thus the assertion $\sum_{k=1}^{\infty}\left\|{\bf w}^{k+1}-{\bf w}^{k}\right\|<+\infty$ holds trivially. Otherwise, since $\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k})$ is nonincreasing from Lemma 1, we have $\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k})>\Theta_{\delta_{1},\delta_{2}}({\bf w}^{*})$ for all $k$. Again from $\lim_{k\rightarrow\infty}\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k})=\Theta_{\delta_{1},\delta_{2}}({\bf w}^{*})$, we know that for any $\eta>0$, there exists a nonnegative integer $k_{0}$ such that $\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k})<\Theta_{\delta_{1},\delta_{2}}({\bf w}^{*})+\eta$ for any $k>k_{0}$. In addition, for any $\varsigma>0$ there exists a positive integer $k_{1}$ such that ${\rm dist}({\bf w}^{k},z({\bf w}^{0}))<\varsigma$ for all $k>k_{1}$. Consequently, for any $\eta,~{}\varsigma>0$, when $k>k_{2}:=\max\{k_{0},k_{1}\}$, we have
${\rm dist}({\bf w}^{k},z({\bf w}^{0}))<\varsigma\quad\hbox{and}\quad\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k})<\Theta_{\delta_{1},\delta_{2}}({\bf w}^{*})+\eta.$
Since $z({\bf w}^{0})$ is a nonempty and compact set, and $\Theta_{\delta_{1},\delta_{2}}$ is a constant on $z({\bf w}^{0})$, we can apply Lemma 4 with $\Omega:=z({\bf w}^{0})$. Therefore, for any $k>k_{2}$, we have
${\varphi_{k}}^{\prime}(\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k})-\Theta_{\delta_{1},\delta_{2}}({\bf w}^{*})){\rm dist}(0,\partial\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k}))\geq 1.$ (45)
From the concavity of ${\varphi_{k}}$, we have
${\varphi_{k}}(\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k})-\Theta_{\delta_{1},\delta_{2}}({\bf w}^{*}))-{\varphi_{k}}(\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k+1})-\Theta_{\delta_{1},\delta_{2}}({\bf w}^{*}))$
$\geq{\varphi_{k}}^{\prime}(\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k})-\Theta_{\delta_{1},\delta_{2}}({\bf w}^{*}))(\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k})-\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k+1})).$
For convenience, for all $p,q\in\mathbb{N}$, we define
$\zeta_{p,q}:=$ ${\varphi_{k}}\big{(}\Theta_{\delta_{1},\delta_{2}}(u^{p})-\Theta_{\delta_{1},\delta_{2}}({\bf w}^{*})\big{)}$
$-{\varphi_{k}}\big{(}\Theta_{\delta_{1},\delta_{2}}(u^{q})-\Theta_{\delta_{1},\delta_{2}}({\bf w}^{*})\big{)}.$
[…]
$\begin{split}&\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k})-\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k+1})\\ &\leq\frac{\zeta_{k,k+1}}{{\varphi_{k}}^{\prime}(\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k})-\Theta_{\delta_{1},\delta_{2}}({\bf w}^{*}))}\\ &\leq\gamma\|{\bf w}^{k}-{\bf w}^{k-1}\|\zeta_{k,k+1}.\end{split}$
Combining Lemma 1 and the above relation, it yields that for any $k>k_{2}$,
$\delta\left\|\mathbf{w}^{k+1}-\mathbf{w}^{k}\right\|^{2}$
$\leq\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k})-\Theta_{\delta_{1},\delta_{2}}({\bf w}^{k+1})$
$\leq\gamma\|{\bf w}^{k}-{\bf w}^{k-1}\|\zeta_{k,k+1},$
This implies that
$\left\|\mathbf{w}^{k+1}-\mathbf{w}^{k}\right\|$
$\leq\sqrt{\frac{\gamma}{\delta}\zeta_{k,k+1}\|{\bf w}^{k}-{\bf w}^{k-1}\|}$
$\leq\frac{\gamma}{2\delta}\zeta_{k,k+1}+\frac{1}{2}\|{\bf w}^{k}-{\bf w}^{k-1}\|,$
which implies that $\sum_{k=1}^{\infty}\left\|{\bf w}^{k+1}-{\bf w}^{k}\right\|<+\infty$. Thus, $\{{\bf w}^{k}\}_{k\geq 1}$ is a Cauchy sequence and hence it is convergent, which also implies that $\{(x_{1}^{k},x_{2}^{k})\}_{k\in\mathbb{N}}$ is convergent. Suppose that $\lim_{k\rightarrow\infty}(x_{1}^{k},x_{2}^{k})=(x_{1}^{*},x_{2}^{*})$. Then, we give that $0\in\partial\Phi(x_{1}^{*},x_{2}^{*})$. Hence, $w_{1}^{*}=\left(w_{11}^{*},w_{12}^{*}\right)=(x_{1}^{*},x_{2}^{*}),w_{2}^{*}=\left(w_{21}^{*},w_{22}^{*}\right)=(x_{1}^{*},x_{2}^{*})$, and $\mathbf{w}^{*}=\left(w_{1}^{*},w_{2}^{*}\right)$. $0\in\Theta_{\delta_{1},\delta_{2}}({\bf w}^{*})$ implies that
0( 0$\in$(0 ∈ ( $\partial_{w_{11}}\Phi(w_{1}^{*})+\delta_{1}(w_{11}^{*}-w_{12}^{*}),\partial_{w_{12}}\Phi(w_{1}^{*})+\delta_{2}(w_{21}^{*}-w_{22}^{*}),$
$\delta_{1}(w_{12}^{*}-w_{11}^{*}),\delta_{2}(w_{22}^{*}-w_{21}^{*})),$
which means that
$0\in(\partial_{w_{11}}\Phi(w_{1}^{*}),\partial_{w_{12}}\Phi(w_{1}^{*}))=\partial\Phi(x_{1}^{*},x_{2}^{*}).$
Therefore, $(x_{1}^{*},x_{2}^{*})$ is a critical point of (13). This completes the proof. ∎